OpenStudy (anonymous):

How do I find the principal angle of 500 degrees?

OpenStudy (amistre64):

divide it by 360 im sure

OpenStudy (anonymous):

by going out to look for it

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

For example, the principal value of -90degrees is 270 degrees.

OpenStudy (amistre64):

500/360 = 50/36 = 25/18 = 1 with a remainder of 7 its 360/7 i think

OpenStudy (anonymous):

thats strange, in the solutions paper. I get -220

OpenStudy (amistre64):

500 -360 ----- 140 thatsa it

OpenStudy (anonymous):

In the solutions paper, the principal angle of 500degrees is -220degrees though

OpenStudy (amistre64):

whats the definition of a principal angle?

OpenStudy (anonymous):

the angle between 0degrees and 360degrees

OpenStudy (anonymous):

so the principal angle of -90 is 270

OpenStudy (anonymous):

that makes sences, But im not sure how to do it with larger numbers like 500

OpenStudy (anonymous):

Thats the example

OpenStudy (amistre64):

360 + principal angle = 500 principal angle = 500 - 360

OpenStudy (amistre64):

might be a scalar for larger number tho but since 500 is less the twice 360 its a safe bet

OpenStudy (anonymous):

Ohh. but if 500-360 gives me 140degrees, why in the solution page, i get -220?

OpenStudy (anonymous):

How can I get -220 as a reult?

OpenStudy (amistre64):

im not all that concerned about a "solution" page

OpenStudy (anonymous):

ok, so the principal angle of 1250 is -190

OpenStudy (anonymous):

how do i get that?

OpenStudy (amistre64):